This 4 pts I3 of 20 (9 complete) ▼ Excess reserves C A. are reserves banks keep to meet the reserve requirement are loans made at above market interest rates. C. are reserves banks keep above the legal requirement O D. are the deposits that banks do not use to make loans. Suppose the required reserve ratio is 12% and a bank has the following balance sheet Assets Reserves $3.000 Deposits $15,000 Loans $12,000 This bank keeps required reserves of S□ and excess reserves of Se Enter your responsesas negers)

Hello can some one help me calculate this problem
Answer #1

Excess reserve means an excess of what is required by bank, regulators, creditors. Excess reserve= Reserve-required reserve. Therefore the option C is correct.

Required reserve= required reserve ratio* demand deposit=12%*15000=0.12*15000=$1800

Excess reserve= Reserve-required reserve=3000-1800=$1200.